# Entanglement and particle production from cosmological perturbations: a quantum optical simulation approach

### v2

@@ 1 Introduction
An interesting property of the two-mode squeezed state of the cosmological perturbations is that it belongs to the class of Gaussian states in continuous-variable (CV) quantum optics. Its Wigner function is Gaussian, and its entanglement and correlations are fully characterized by its covariance matrix Weedbrook_2012 ; ferraro2005gaussianstatescontinuousvariable ; serafini2017quantum_primer ; Adhikari:2022oxr ; Adhikari:2022whf ; Adhikari:2023evu ; Li:2023ekd ; Li:2024kfm ; Rabinovici:2025otw . This parallelism makes it possible to apply methods created for optical systems, especially those that use continuous-variable Gaussian states, directly to inflationary cosmology. Each wave vector $k$ and the pair of modes (k,k)(k,-k) of cosmological perturbations evolve under a linear quadratic Hamiltonian, which evolves the quantum state through a symplectic transformation acting on the covariance matrix. Importantly, any two-mode covariance matrix can be diagonalized using a symplectic transformation. The symplectic eigenvalues of the diagonalized covariance matrix encode the quantum uncertainties and the entanglement of the state ferraro2005gaussianstatescontinuousvariable ; plenio2007_entanglement_measures .
Likewise, non-Gaussian features in cosmological perturbations are naturally produced by non-linear gravitational effects, indicating that the universe is not completely Gaussian. This arises, for instance, in models with a single matter field when taking into account the full cubic action for density perturbations. In simple cases with a canonical scalar field, the speed of sound is $c_{s}=1$; however, in more general models, a reduced $c_{s}$ amplifies equilateral-type non-Gaussianity and strengthens cubic interactions brahma2020entanglement . Such cubic terms modify the Hamiltonian beyond quadratic form, even in minimal models, making it increasingly difficult to analytically solve these non-Gaussian features classically. However, quantum optics platforms offer a natural toolkit for simulating non-Gaussian dynamics of cosmological perturbations using cubic phase gates in inflationary cosmology.
It is a widely accepted fact that the inflationary cosmology paradigm is the ‘standard model’ of the hot big bang cosmology guth1987inflationary . Inflation not only solves the horizon and the flatness problem but also predicts the sets of perturbations on the Friedmann-Lemaitre-Robertson-Walker (FLRW) background metric. These perturbations are the results of the amplification of quantum fluctuations, during inflation, into macroscopic perturbations which later evolve into large-scale structures such as galaxies, clusters, and cosmic microwave background anisotropies mukhanov1982vacuum .
@@
These small perturbations, $\delta g_{\mu\nu}$ and $\delta\varphi(\mathbf{x},\eta)$, are treated quantum mechanically while the background obeys the classical Friedmann equation mukhanov1992theory ; straumann2006primordial ; dodelson2024modern . It is sensible to treat the metric perturbation itself quantum mechanically during inflation, as the energy density during this epoch is so enormous that we expect quantum-gravity effects to play a role.
For the purpose of studying cosmological perturbations, it turns out that both the scalar metric and the field fluctuations are encoded in a single degree of freedom, the comoving curvature perturbation $\mathcal{R}$, which is a gauge-invariant variable. In Refs. mukhanov1985gravitational ; sasaki1986large , Mukhanov and Sasaki introduced a variable $v$, that is defined as:
@@
Moreover, the quantization of these perturbations during inflation turns their vacuum fluctuations into real particles. By canonically quantizing these modes and by imposing a Bunch-Davies vacuum state, so that each mode matches a positive-frequency plane wave for $k\gg aH$ (sub-Hubble region), the subsequent evolution mixes positive and negative frequencies. This mode mixing is most easily described by Bogoliubov transformations, where the early time annihilation operators $\hat{c}_{\mathbf{k}}$ relate to the late time operators $b_{\mathbf{k}}$ via
@@
with $|\alpha_{k}|^{2}-|\beta_{k}|^{2}=1$. A nonzero $\beta_{k}$ means that what started out as a vacuum state at early times ends up containing particles when viewed at late times, with each mode having an average particle number of $|\beta_{k}|^{2}$ mukhanov1992theory ; parker1969quantized ; jiang2025cosmological .
@@
An alternative, yet equivalent, description of particle production is given through squeezed quantum states. In mukhanov1992theory ; albrecht1994inflation it was shown that the amplification of ground state quantum fluctuations during inflation is a process of quantum squeezing. Within sub-Hubble scales, the Fourier modes of these perturbations undergo oscillations. However, during the accelerated expansion of inflation, the physical wavelengths of these modes grow faster than the Hubble radius $H^{-1}(t)$. As a result, the modes eventually exit the Hubble horizon. Once they cross this threshold, their quantum state evolves into a squeezed vacuum state grishchuk1990squeezed ; prokopec1994classical . The squeezing of quantum field modes during inflation produces particle pairs with opposing momenta via amplified quantum fluctuations. Hence, the cosmological perturbation in the FLRW background metric can be described in the language of a two-mode squeezed state.
@@
The hallmark of the two-mode squeezed state is that the perturbed modes $k$ and k-k are highly entangled. The strong correlations between the entangled modes imply that the reduced density matrix of the squeezed state, which is mixed through a subtle coarse-graining, gives the non-vanishing von Neumann entropy. In brahma2020entanglement the coarse-graining entropy was done by averaging over the squeezing angle, which provides a stochastic part that lead the reduced density matrix’s off diagonal elements to zero. This process gives rise to entanglement entropy, which can account for part of the large entropy observed in the post-inflationary universe gasperini1993entropy ; boutivas2023entanglement ; martin2021real ; Martin:2021xml . Also, many of the papers have argued that the squeezing phenomenon in the context of cosmology is associated with the emergence of classical behavior of our early quantum universe grishchuk1990squeezed ; albrecht1994inflation ; polarski1996semiclassicality ; lesgourgues1997quantum ; kiefer1998quantum ; kiefer2009quantum ; perez2006quantum ; brandenberger1990classical ; calzetta1995quantum ; Li:2024ljz ; Li:2024iji ; colas2022four .
@@
This work contributes to the growing efforts of applying quantum optics and quantum information theory to the study of inflationary cosmology. Recent studies have demonstrated that quantum optical methods like continuous-variable quantum circuits, Gaussian-state manipulation, and entanglement measures provide powerful frameworks to explore primordial quantum states Liu_2021 ; Boutivas_2024 . Several works have explored simulating inflationary cosmology using quantum tools. For example, digital simulation of particle creation in expanding universes were done using quantum circuits and extended field theory algorithms Lloyd:2013xba ; Bao:2017iye ; Yang_2020 ; Maceda_2025 ; Piotrak:2025zhy . The redistribution of Gaussian entanglement for continuous variable of scalar field in an expanding universe was also analyzed by modeling the expansion as a Gaussian channel Li2022GaussianEntanglement . Earlier studies showed that quantum fluctuations during inflation naturally form squeezed states parker1969quantized ; grishchuk1990quantum . Continuous variable photonic systems have been proposed to simulate quantum fields using the Gaussian formalism Marshall_2015 . Other works studied entropy and decoherence, linking squeezing to the emergence of classical structures grishchuk1990quantum ; albrecht1994inflationary ; Zhai:2024odw . Together, these works provide valuable context and motivation for our approach.
@@
In addition to quantum-optical simulations, condensed-matter systems have long been studied as analog models of cosmology. Analog models like Bose-Einstein condensates and coupled mechanical oscillators have been used as laboratory analogues of the early universe mimicking inflation-like expansion and particle pair production. Fischer_2004 ; Rhyno_2025 ; PhysRevA.103.023322 ; PhysRevD.110.125023 ; PhysRevLett.123.180502 . Experiment with the quantum fluids of light have been shown that a sudden change in a 33D photon fluid can cause spontaneous creation of particles, forming acoustic-peak patterns like those cosmic inflation Steinhauer_2022 . Various systems from expanding superfluid rings to photonic condensates and optical fiber circuits have been explored as analogues for early universe studies Pal_2024 . Other works include probing sonic analogues of Gibbons–Hawking radiation, trans-Planckian effects, and quantum backreaction, and exploring phenomena such as analogue gravitational lensing, black hole information scrambling, nonlinear quantum effects, and phase transitions in dipolar systems Fedichev2003PRL ; Fischer2004PRA ; Fischer2022EPJC ; Fischer2022PRD ; Fedichev2004PRD ; Fischer2017PRL ; Fischer2022PRA ; Fischer2023CQG ; Fischer2023PRD ; Fischer2024PRD ; Fischer2025arxiv ; Visser2002PRL . Besides quantum-optical and condensed-matter settings, synthetic mechanical lattices have also been employed to study analogues of inflationary and FLRW cosmology. In Refs. rhyno2025mechanical it is demonstrated how the bosonic dynamics of scalar-field fluctuations are mapped onto the motion of mechanical oscillators, and key inflationary quantities such as particle production and the power spectrum are simulated and measured. These above approaches from the divergent field in contrast to circuit based method we employed, creates the motivation to explore more on this subject matter and discover the unknown.
@@
In this work, we use tools and techniques from quantum optics to simulate the inflationary squeezing Hamiltonian. We use Gaussian states and unitaries to model a two-mode squeezed vacuum and draw a quantum optics circuit Gerry_Knight_2004 . We use this circuit to compute entanglement measures like von Neumann entropy and logarithmic negativity of the system. We also analyze the effect of thermal noise on the quantum state and different entanglement measures plenio2007_entanglement_measures ; Gerry_Knight_2004 ; Plenio_entanglement ; Vidal_Werner_entanglement ; Choudhury:2016cso ; Choudhury:2016pfr ; Brady_Agullo_Kranas_2022 ; kranas2023entanglement . Moreover, we simulate the entropy measures of cosmological perturbations across various cosmological backgrounds. In particular, we focus on the behavior of the entanglement entropy for backgrounds with an arbitrary, constant equation of state $w$, covering both expanding (accelerating and decelerating) and contracting (accelerating and decelerating) backgrounds.
@@ 2.1 Review of the formalism of cosmological perturbations
We adopt the gauge-invariant approach for the equation of motion of the metric and scalar field perturbation. It ensures physical predictions are free from dependency on coordinate transformations mukhanov1992theory . We, therefore, first expand the FLRW metric around a background state and introduce scalar perturbations with four degrees of freedom. We simplify the metric by choosing an appropriate gauge, such as the longitudinal gauge, which reduces the number of independent variables. Then the total action consisting of both gravitational and matter components is minimized to express the equation of motion of the perturbation of the scalar field in terms of a gauge-invariant quantity, which is, in our case, the Mukhanov-Sasaki variable.
@@ 2.1.1 Longitudinal gauge invariant metric
There are four degrees of freedom in $\delta g_{\mu\nu}$ which correspond to scalar metric fluctuations. In cosmology, perturbations can be scalar, vector, or tensor. Here, we focus on scalar perturbations (related to density and pressure fluctuations), which have four independent scalar functions.
@@
which is a gauge artifact and carries no physical significance brandenberger2004lectures .
@@ 2.1.2 Evolution equations
We utilize the Mukhanov-Sasaki formalism to derive the equation of motion of the scalar perturbation mukhanov1992theory . The total action is given as:
@@
Expanding about the background field $\varphi_{0}$ using a Taylor series, we get the first perturbation of the Lagrangian and the second perturbation of the Lagrangian, where $L_{0}$ is the Lagrangian density evaluated at the background metric and $g_{0}$ is the background metric. Now, applying the background equations. 20, 21, 22 and integrating by parts and combining the second-order variations of the matter action $\delta_{2}S_{m}$ and the Einstein-Hilbert action $\delta_{2}S_{EH}$ (17) and utilizing the longitudinal gauge, we obtain
@@
This equation relates the perturbations $\phi$ to the perturbation of the scalar field $\delta\varphi$. However, these variables are gauge-dependent, meaning their values depend on the choice of coordinate system. To eliminate this dependence, we define the gauge-invariant Mukhanov-Sasaki variable mukhanov1985gravitational ; sasaki1986large :
@@
where, $\epsilon=-\frac{\dot{H}}{H^{2}}$, that is first slow-roll inflation parameter. After transforming the action into terms of the gauge-invariant variable v, the second-order perturbed action mukhanov1992theory simplifies to:
@@ 2.2 Quantization and formalism of squeezed states
The classical analysis of the fluctuation is sufficient to describe how the perturbations evolve. However, to study the origin of the perturbations and particle production in expanding background space times quantum treatment is necessary brandenberger1993classical . In this section, we investigate the quantum nature of the perturbations to see how the expanding background leads to particle production. We treat them as field that evolves in an expanding spacetime. Using the Heisenberg picture, we construct the Hamiltonian and the equation of motion of ladder operators by using the Heisenberg equation of motion. And we do the Bogoliubov transformation to see the evolution of the field modes from the initial time $\eta_{0}$ (vacuum).

### v3

@@ 1 Introduction
An interesting property of the two-mode squeezed state of the cosmological perturbations is that it belongs to the class of Gaussian states in continuous-variable (CV) quantum optics. Its Wigner function is Gaussian, and its entanglement and correlations are fully characterized by its covariance matrix Ferraro:2005hen ; Weedbrook:2011wxo ; Serafini:2017rrn ; Adhikari:2022oxr ; Adhikari:2022whf ; Adhikari:2023evu ; Li:2023ekd ; Li:2024kfm ; Rabinovici:2025otw . This parallelism makes it possible to apply methods created for optical systems, especially those that use continuous-variable Gaussian states, directly to inflationary cosmology. Each wave vector $k$ and the pair of modes (k,k)(k,-k) of cosmological perturbations evolve under a linear quadratic Hamiltonian, which evolves the quantum state through a symplectic transformation acting on the covariance matrix. Importantly, any two-mode covariance matrix can be diagonalized using a symplectic transformation. The symplectic eigenvalues of the diagonalized covariance matrix encode the quantum uncertainties and the entanglement of the state Ferraro:2005hen ; Plenio:2007zz .
Likewise, non-Gaussian features in cosmological perturbations are naturally produced by non-linear gravitational effects, indicating that the universe is not completely Gaussian. This arises, for instance, in models with a single matter field when taking into account the full cubic action for density perturbations. In simple cases with a canonical scalar field, the speed of sound is $c_{s}=1$; however, in more general models, a reduced $c_{s}$ amplifies equilateral-type non-Gaussianity and strengthens cubic interactions Brahma:2020zpk . Such cubic terms modify the Hamiltonian beyond quadratic form, even in minimal models, making it increasingly difficult to analytically solve these non-Gaussian features classically. However, quantum optics platforms offer a natural toolkit for simulating non‑Gaussian dynamics of cosmological perturbations using cubic phase gates in inflationary cosmology.
It is a widely accepted fact that the inflationary cosmology paradigm is the ‘standard model’ of the hot big bang cosmology Guth:1980zm . Inflation not only solves the horizon and the flatness problem but also predicts the sets of perturbations on the Friedmann-Lemaitre-Robertson-Walker (FLRW) background metric. These perturbations are the results of the amplification of quantum fluctuations, during inflation, into macroscopic perturbations which later evolve into large-scale structures such as galaxies, clusters, and cosmic microwave background anisotropies Mukhanov:1982nu .
@@
These small perturbations, $\delta g_{\mu\nu}$ and $\delta\varphi(\mathbf{x},\eta)$, are treated quantum mechanically while the background obeys the classical Friedmann equation Mukhanov:1990me ; Dodelson:2003ft ; Straumann:2005mz . It is sensible to treat the metric perturbation itself quantum mechanically during inflation, as the energy density during this epoch is so enormous that we expect quantum-gravity effects to play a role.
For the purpose of studying cosmological perturbations, it turns out that both the scalar metric and the field fluctuations are encoded in a single degree of freedom, the comoving curvature perturbation $\mathcal{R}$, which is a gauge-invariant variable. In Refs. Mukhanov:1985rz ; Sasaki:1986hm , Mukhanov and Sasaki introduced a variable $v$, that is defined as:
@@
Moreover, the quantization of these perturbations during inflation turns their vacuum fluctuations into real particles. By canonically quantizing these modes and by imposing a Bunch-Davies vacuum state, so that each mode matches a positive‐frequency plane wave for $k\gg aH$ (sub-Hubble region), the subsequent evolution mixes positive and negative frequencies. This mode mixing is most easily described by Bogoliubov transformations, where the early time annihilation operators $\hat{c}_{\mathbf{k}}$ relate to the late time operators $b_{\mathbf{k}}$ via
@@
with $|\alpha_{k}|^{2}-|\beta_{k}|^{2}=1$. A nonzero $\beta_{k}$ means that what started out as a vacuum state at early times ends up containing particles when viewed at late times, with each mode having an average particle number of $|\beta_{k}|^{2}$ Parker:1969au ; Mukhanov:1990me ; Jiang:2025ktt .
@@
An alternative, yet equivalent, description of particle production is given through squeezed quantum states. In Mukhanov:1990me ; Albrecht:1992kf it was shown that the amplification of ground state quantum fluctuations during inflation is a process of quantum squeezing. Within sub-Hubble scales, the Fourier modes of these perturbations undergo oscillations. However, during the accelerated expansion of inflation, the physical wavelengths of these modes grow faster than the Hubble radius $H^{-1}(t)$. As a result, the modes eventually exit the Hubble horizon. Once they cross this threshold, their quantum state evolves into a squeezed vacuum state Grishchuk:1990bj ; Albrecht:1992kf . The squeezing of quantum field modes during inflation produces particle pairs with opposing momenta via amplified quantum fluctuations. Hence, the cosmological perturbation in the FLRW background metric can be described in the language of a two-mode squeezed state.
@@
The hallmark of the two-mode squeezed state is that the perturbed modes $k$ and k-k are highly entangled. The strong correlations between the entangled modes imply that the reduced density matrix of the squeezed state, which is mixed through a subtle coarse-graining, gives the non-vanishing von Neumann entropy. In Brahma:2020zpk the coarse-graining entropy was done by averaging over the squeezing angle, which provides a stochastic part that lead the reduced density matrix’s off diagonal elements to zero. This process gives rise to entanglement entropy, which can account for part of the large entropy observed in the post-inflationary universe Gasperini:1992xv ; Martin:2021qkg ; Martin:2021xml ; Boutivas:2023ksg . Also, many of the papers have argued that the squeezing phenomenon in the context of cosmology is associated with the emergence of classical behavior of our early quantum universe Brandenberger:1990bx ; Grishchuk:1990bj ; Albrecht:1992kf ; Polarski:1995jg ; Lesgourgues:1996jc ; Kiefer:1998qe ; Calzetta:1995ys ; Perez:2005gh ; Kiefer:2008ku ; Colas:2021llj ; Li:2024ljz ; Li:2024iji .
@@
This work contributes to the growing efforts of applying quantum optics and quantum information theory to the study of inflationary cosmology. Recent studies have demonstrated that quantum optical methods like continuous-variable quantum circuits, Gaussian-state manipulation, and entanglement measures provide powerful frameworks to explore primordial quantum states Liu:2020wtr ; Boutivas:2023mfg . Several works have explored simulating inflationary cosmology using quantum tools. For example, digital simulation of particle creation in expanding universes were done using quantum circuits and extended field theory algorithms Lloyd:2013xba ; Bao:2017iye ; Yang:2019kbb ; Maceda:2024rrd ; Piotrak:2025zhy . The redistribution of Gaussian entanglement for continuous variable of scalar field in an expanding universe was also analyzed by modeling the expansion as a Gaussian channel Li:2023gtf . Earlier studies showed that quantum fluctuations during inflation naturally form squeezed states Parker:1969au ; Grishchuk:1993ds . Continuous variable photonic systems have been proposed to simulate quantum fields using the Gaussian formalism Marshall:2015mna . Other works studied entropy and decoherence, linking squeezing to the emergence of classical structures Albrecht:1992kf ; Grishchuk:1993ds ; Zhai:2024odw . Together, these works provide valuable context and motivation for our approach.
@@
In addition to quantum-optical simulations, condensed-matter systems have long been studied as analog models of cosmology. Analog models like Bose-Einstein condensates and coupled mechanical oscillators have been used as laboratory analogues of the early universe mimicking inflation-like expansion and particle pair production Fischer:2004bf ; Wittemer:2019agm ; Bhardwaj:2020ndh ; Rhyno:2023kud ; Agullo:2024lry . Experiment with the quantum fluids of light have been shown that a sudden change in a 33D photon fluid can cause spontaneous creation of particles, forming acoustic-peak patterns like those cosmic inflation Steinhauer:2021fhb . Various systems from expanding superfluid rings to photonic condensates and optical fiber circuits have been explored as analogues for early universe studies Pal:2024qno . Other works include probing sonic analogues of Gibbons–Hawking radiation, trans-Planckian effects, and quantum backreaction, and exploring phenomena such as analogue gravitational lensing, black hole information scrambling, nonlinear quantum effects, and phase transitions in dipolar systems Fischer:2001jz ; Fedichev:2003id ; Fedichev:2003bv ; Fedichev:2003dj ; Cha:2016esj ; Tian:2020bze ; Ribeiro:2021fpk ; Ribeiro:2022gln ; Baak:2022hum ; Baak:2023zjf ; Pal:2024qno ; Chandran:2025azu . Besides quantum-optical and condensed-matter settings, synthetic mechanical lattices have also been employed to study analogues of inflationary and FLRW cosmology. In Refs. Rhyno:2023kud it is demonstrated how the bosonic dynamics of scalar-field fluctuations are mapped onto the motion of mechanical oscillators, and key inflationary quantities such as particle production and the power spectrum are simulated and measured. These above approaches from the divergent field in contrast to circuit based method we employed, creates the motivation to explore more on this subject matter and discover the unknown.
@@
In this work, we use tools and techniques from quantum optics to simulate the inflationary squeezing Hamiltonian. We use Gaussian states and unitaries to model a two-mode squeezed vacuum and draw a quantum optics circuit Gerry_Knight_2004 . We use this circuit to compute entanglement measures like von Neumann entropy and logarithmic negativity of the system. We also analyze the effect of thermal noise on the quantum state and different entanglement measures Vidal:2002zz ; Gerry_Knight_2004 ; Plenio:2005cwa ; Plenio:2007zz ; Choudhury:2016cso ; Choudhury:2016pfr ; Brady:2022ffk ; Kranas:2023aph . Moreover, we simulate the entropy measures of cosmological perturbations across various cosmological backgrounds. In particular, we focus on the behavior of the entanglement entropy for backgrounds with an arbitrary, constant equation of state $w$, covering both expanding (accelerating and decelerating) and contracting (accelerating and decelerating) backgrounds.
@@ 2.1 Review of the formalism of cosmological perturbations
We adopt the gauge-invariant approach for the equation of motion of the metric and scalar field perturbation. It ensures physical predictions are free from dependency on coordinate transformations Mukhanov:1990me . We, therefore, first expand the FLRW metric around a background state and introduce scalar perturbations with four degrees of freedom. We simplify the metric by choosing an appropriate gauge, such as the longitudinal gauge, which reduces the number of independent variables. Then the total action consisting of both gravitational and matter components is minimized to express the equation of motion of the perturbation of the scalar field in terms of a gauge-invariant quantity, which is, in our case, the Mukhanov-Sasaki variable.
@@ 2.1.1 Longitudinal gauge invariant metric
There are four degrees of freedom in $\delta g_{\mu\nu}$ which correspond to scalar metric fluctuations. In cosmology, perturbations can be scalar, vector, or tensor. Here, we focus on scalar perturbations (related to density and pressure fluctuations), which have four independent scalar functions
@@
which is a gauge artifact and carries no physical significance Brandenberger:2003vk .
@@ 2.1.2 Evolution equations
We utilize the Mukhanov-Sasaki formalism to derive the equation of motion of the scalar perturbation Mukhanov:1990me . The total action is given as:
@@
Expanding about the background field $\varphi_{0}$ using a Taylor series, we get the first perturbation of the Lagrangian and the second perturbation of the Lagrangian, where $L_{0}$ is the Lagrangian density evaluated at the background metric and $g_{0}$ is the background metric. Now, applying the background equations 20, 21, 22 and integrating by parts and combining the second-order variations of the matter action $\delta_{2}S_{m}$ and the Einstein-Hilbert action $\delta_{2}S_{EH}$ (17) and utilizing the longitudinal gauge, we obtain
@@
This equation relates the perturbations $\phi$ to the perturbation of the scalar field $\delta\varphi$. However, these variables are gauge-dependent, meaning their values depend on the choice of coordinate system. To eliminate this dependence, we define the gauge-invariant Mukhanov-Sasaki variable Mukhanov:1985rz ; Sasaki:1986hm :
@@
where, $\epsilon=-\frac{\dot{H}}{H^{2}}$, that is first slow-roll inflation parameter. After transforming the action into terms of the gauge-invariant variable v, the second-order perturbed action Mukhanov:1990me simplifies to:
@@ 2.2 Quantization and formalism of squeezed states
The classical analysis of the fluctuation is sufficient to describe how the perturbations evolve. However, to study the origin of the perturbations and particle production in expanding background space times quantum treatment is necessary Brandenberger:1993zc . In this section, we investigate the quantum nature of the perturbations to see how the expanding background leads to particle production. We treat them as field that evolves in an expanding spacetime. Using the Heisenberg picture, we construct the Hamiltonian and the equation of motion of ladder operators by using the Heisenberg equation of motion. And we do the Bogoliubov transformation to see the evolution of the field modes from the initial time $\eta_{0}$ (vacuum).